Solve the open sentence. then graph the solution set.
l r+3 l = 10 l=absolute value marks
i didnt get this at all in class today, please help me.

Answer by stanbon(75887)

| r+3 | = 10
If the left side is 10 the r+3 must be 10 or it must be -10
So,
If r+3=10
then r=7.
But, if r+3=-10
then r=-13
-----------------
So you get two solutions: r=7 or r=-13
-------
Cheers,
Stan H.